# On the gravitational stratification of multi-fluid-multi-species plasma

### 17 Mar 2026 v2

@@ 4 MFMS equations and numerical methods
In the following, we run 1D MFMS numerical simulations that use the stratifications considered above as initial conditions. We use the numerical code Ebysus (Martínez-Sykora et al., 2020b, 2023; Wargnier et al., 2023, 2025). Compared to various multi-fluid codes (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Popescu Braileanu & Keppens, 2022; Snow et al., 2023), an important feature of Ebysus is that it can solve the MHD equations separately for each of the desired number of excited levels, ionization stages and species, although in this work, we solve the collective motion of all the excited levels of each neutral species for simplicity. Ebysus is modular and allows for the inclusion of ionization/excitation in NEQ, collision between different fluids, thermal conduction, radiative losses, and the Hall term. Details of the derivation of the MFMS formulae have been introduced by Khomenko et al. (2014); Ballester et al. (2018). For the purpose of this work, we neglect the Hall term and thermal conduction, and consider only 1D.
@@
To spatially discretize the governing equations, Ebysus uses a sixth-order finite-difference scheme with hyper-diffusivity. Fifth-order interpolation is used to interpolate variables in the staggered grid system. To advance in time, we use the modified explicit third-order predictor-corrector Hyman method (Hyman, 1979). These numerical approaches are inherited from the Bifrost code (Gudiksen et al., 2011). Recently, we have included in Ebysus the second-order Partitioned Implicit-Explicit Orthogonal Runge-Kutta (PIROCK) method (Abdulle & Vilmart, 2013; Wargnier et al., 2025), allowing for advancing in time by combining efficient explicit stabilized and implicit integration techniques while employing variable time-stepping with error control. PIROCK method is also tested in this work, further supporting the present results.
@@ 5.1 Initial conditions
A cHE stratification obtained using the numerical integration routine is approximately in equilibrium. Therefore, there may be remnant disturbances when a simulation uses the cHE stratification as the initial condition. More importantly, according to our assumption, there must be drift velocities between fluids, which are related to the pressure gradients in the model. To evaluate these effects, we simulate disturbances developing in initially quiet stratifications, without imposing any external dynamic driving force. Specifically, we construct a cHE stratification using the Model C7 temperature distribution, yet including only [^5] The Model C7 provides discrete data points. Therefore, the temperature distribution is interpolated using the interp1d function from the NumPy library (Harris et al., 2020) with the quadratic interpolation method. While errors still exist, using the linear method would have caused more significant errors in drift velocities and, eventually, in the total velocity. The error in the interpolated temperature distribution is not observable by the naked eye, but it affects the sound speed and pressure gradient. 2.2 to 3.2 Mm and hydrogen and helium (no He2+). Therefore, the domain does not include the non-smooth transition from the upper chromosphere to the TR at around 2.14 Mm in the Model C7. We include only the smooth (yet steep) variation above the transition, where the pressure gradient is large.
@@ 6 The ponderomotive force and the resulting motions
Since significant decoupling appears in the cHE stratification when the pressure gradient is large, it is interesting to know how this ”static” decoupling would contribute to the MFMS dynamics. Here, we run simulations of a monochromatic Alfvén wave, using the cHE and pHE stratifications introduced in the previous section. Alfvén waves existing in the low solar atmosphere have been suggested (De Pontieu et al., 2007; McIntosh et al., 2011; Chae et al., 2021), and they may play a vital role in coronal heating (Van Doorsselaere et al., 2020), the FIP effect (Laming, 2004, 2015), etc., which are beyond the scope of this work. We are interested in the fact that the ponderomotive force induced by Alfvén waves drives plasma motions along the magnetic field (Lundin & Guglielmi, 2006), and such motions can be simulated and used as an example to compare the effects of using different stratifications.
@@
Figure 8: Longitudinal velocities at $t=1$ s. All settings are listed in Table 1. Except simulation (d), which uses the pHE stratification, all the others use the cHE stratification as the initial condition. Simulations (a) and (b) use assigned constant collisional frequencies. More specifically, (a) is weakly collisional and (b) is strongly collisional, respectively having collisional frequencies $5\times 10^{-2}$ Hz and $10^{2}$ Hz. All the other simulations use physical collisional frequencies calculated by Ebysus. Only simulation (e) includes the NEQ ionization/recombination.
The effects of the numerical disturbances appear in the $z$ direction (i.e., the direction of the magnetic field). The Alfvén wave-induced ponderomotive force relates to ${\partial E^{2}_{\perp}}/{\partial z}$, in which $E_{\perp}$ is the electric field perpendicular to the magnetic field (Lundin & Guglielmi, 2006; Martínez-Sykora et al., 2023). Therefore, upward motions (positive longitudinal velocity) are driven by the ponderomotive force, as shown in Figure 8. When it is weakly collisional (panel a), the neutral fluids are decoupled from the ionized fluids, whose motions are driven by the Alfvén wave. Each neutral fluid also moves independently because of the imbalance between gravity and the respective plasma gas pressure, while neither collisions with ionized fluids nor magnetic pressure affect neutral fluids. In contrast, all fluids in the strongly collisional simulation are coupled, behaving as a single-fluid as expected (panel b). More importantly, we can observe that in the strongly collisional simulation, the total velocity is zero before the wave arrives, showing that the cHE stratification is indeed in hydrostatic equilibrium, providing another way to validate the MFMS code – once it is collisionally coupled, the plasma behaves as a single fluid.
@@
The pHE stratification is not further discussed with ionization and recombination due to the numerical instability. We should, however, note that the ionization and recombination timescales are long in the chromosphere (Carlsson & Stein, 2002), and thus the changes in density and temperature would be less drastic when the pHE stratification is used under chromospheric conditions. However, this also means that the ionization fractions (and thus the collisional interactions) in the corresponding simulations would be largely dependent on the initial stratification. We have shown in Section 2, that heavier elements, whose scale heights are different from that of (ionized or neutral) hydrogen, cannot be properly recovered by the pHE stratification, even if the collisional interactions (or other mixing processes) are sufficient to couple the heavy fluids.
@@ Acknowledgements.
FZ thanks Nicolas Poirier, Mats Carlsson, Boris Gudiksen, and Ben Snow for the helpful discussion and Mikolaj Szydlarski for technical support. This research has been supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622, and by computational resources provided by Sigma2, the National Infrastructure for High Performance Computing and Data Storage in Norway. JMS acknowledges support by NASA contracts NNG09FA40C (IRIS) and 80GSFC21C0011 (MUSE), NASA grant 80NSSC26K0018, NSF grants AGS2532363 and AGS2532187. This work used the CHIANTI atomic database. CHIANTI is a collaborative project involving George Mason University, the University of Michigan (USA), the University of Cambridge (UK), and NASA Goddard Space Flight Center (USA). Figures in this work were produced using the Python package Matplotlib (Hunter, 2007). Data processing was performed using the NumPy library (Harris et al., 2020) and in-house packages EbysusRunTools_py, atom_py and Helita.
@@ Appendix A Effects of magnetic field
Figure 9: The maximum and averaged $|\mathbf{B}|$ at each height. The maximum value shows a clear inverse power law above $\sim$ 0.1 Mm. The averaged value, however, first drastically increases and then gradually decreases, and finally exhibits slow variation above $\sim$1 Mm.
@@
Therefore, following the results in Section 3.3, we discuss the effects of adding a magnetic field. The magnetic field in this 2D model is locally concentrated in the photosphere and shows a significant expansion higher up. Consequently, the unsigned mean magnetic field strength decreases over height (Figure 3, bottom). Observing Figure 9, a power law can be seen from the variation of the maximum value at each height, but it is not representative of the whole magnetic field. The averaged value also cannot represent the derivative (i.e., gradient) of the magnetic field. Nevertheless, we do not intend to focus on such details, and the results justify using an expansion assumption to approximate the effect of magnetic pressure.
@@
where $\mathbf{B}_{0}$ and $r_{0}$ are free parameters, to describe the magnetic pressure from the bottom of the photosphere. With Eq. (15), a smaller $r_{0}$ means that the magnetic field is more concentrated in the photosphere. The pressure gradient can be analytically calculated based on Eq. (15), and directly added to the numerical integration routine introduced in Section 3. With a constant $r_{0}$ and varying $\mathbf{B}_{0}$, we have obtained a series of stratifications with the SE ionization, as shown in Figure 10. Note that only the total densities of all fluids are shown for simplicity. We first acknowledge that without magnetic pressure (gradient), the cHE stratification has a lower density than the averaged 2D solution. In comparison, magnetic pressure gradients significantly alter the solution, as stronger gradients support higher density. More importantly, the results demonstrate that the current numerical integration routine can be used to calculate cHE stratifications under varying magnetic fields.
Figure 10: cHE stratifications calculated with analytical magnetic pressure distributions, $r_{0}=1$ Mm, and SE ionization fractions. Alternatively, NEQ ionization fractions can also be used. The total density when $B_{0}=0$ G is the same as the SE result in Figure 4.
Figure 11: Ratios of the magnetic pressure gradient to the plasma gas pressure gradient shown in Figure 10. The peaks appear because the plasma gas pressure gradient changes sign. Above the photosphere, the magnetic pressure gradient becomes dominant in shaping the stratifications.
@@
in which $L$ is the radius of the loop. This function is essentially the same as what was used by Karampelas et al. (2019). As $z$ is the height from the footpoint, the gravitational acceleration along the loop axis becomes zero at the loop apex ($z=L$). Plasma along a realistic loop is supported by plasma gas pressure and magnetic pressure/tension, and we can already see that the magnetic pressure gradient effectively shapes the 1D stratification. Particularly, Figure 11 shows that the magnetic pressure description given by Eq. (15) results in sufficient magnetic pressure gradients to support the plasma in the upper chromosphere and above. Consequently, including Eq. (16) does not significantly change the stratifications, and the corresponding results are thus not shown.

### 30 Apr 2026 v3

@@ 4 MFMS equations and numerical methods
In the following, we run 1D MFMS numerical simulations that use the stratifications considered above as initial conditions. We use the numerical code Ebysus (Martínez-Sykora et al., 2020b, 2023; Wargnier et al., 2023, 2025). Compared to various multi-fluid codes (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Popescu Braileanu and Keppens, 2022; Snow et al., 2023), an important feature of Ebysus is that it can solve the MHD equations separately for each of the desired number of excited levels, ionization stages and species, although in this work, we solve the collective motion of all the excited levels of each neutral species for simplicity. Ebysus is modular and allows for the inclusion of ionization/excitation in NEQ, collision between different fluids, thermal conduction, radiative losses, and the Hall term. Details of the derivation of the MFMS formulae have been introduced by Khomenko,E. et al. (2014); Ballester et al. (2018). For the purpose of this work, we neglect radiation, thermal conduction, and the Hall term, and consider only 1D.
@@
To spatially discretize the governing equations, Ebysus uses a sixth-order finite-difference scheme with hyper-diffusivity. Fifth-order interpolation is used to interpolate variables in the staggered grid system. To advance in time, we use the modified explicit third-order predictor-corrector Hyman method (Hyman, 1979). These numerical approaches are inherited from the Bifrost code (Gudiksen et al., 2011). Recently, we have included in Ebysus the second-order Partitioned Implicit-Explicit Orthogonal Runge-Kutta (PIROCK) method (Abdulle and Vilmart, 2013; Wargnier et al., 2025), allowing for advancing in time by combining efficient explicit stabilized and implicit integration techniques while employing variable time-stepping with error control. PIROCK method is also tested in this work, further supporting the present results.
@@ 5.1 Initial conditions
A cHE stratification obtained using the numerical integration routine is approximately in equilibrium. Therefore, there may be remnant disturbances when a simulation uses the cHE stratification as the initial condition. More importantly, according to our assumption, there must be drift velocities between fluids. To evaluate these effects, we simulate disturbances developing in initially quiet stratifications, without imposing any external dynamic driving force. Specifically, we construct a cHE stratification using the Model C7 temperature distribution, yet including only [^5] The Model C7 provides discrete data points. Therefore, the temperature distribution is interpolated using the interp1d function from the NumPy library (Harris et al., 2020) with the quadratic interpolation method. Using the linear method would have caused more significant errors in drift velocities and, eventually, in the total velocity. 2.2 to 3.2 Mm and hydrogen and helium (no He2+). Therefore, the domain does not include the non-smooth transition from the upper chromosphere to the TR at around 2.14 Mm in the Model C7. We include only the smooth (yet steep) variation above the transition, where the pressure gradient is large.
@@ 6 The ponderomotive force and the resulting motions
Since significant decoupling appears in the cHE stratification when the pressure gradient is large, it is interesting to know how this ”static” decoupling would contribute to the MFMS dynamics. Here, we run simulations of a monochromatic Alfvén wave, using the cHE and pHE stratifications introduced in the previous section. Alfvén waves existing in the low solar atmosphere have been suggested (De Pontieu et al., 2007; McIntosh et al., 2011; Chae et al., 2021), and they may play a vital role in coronal heating (Van Doorsselaere et al., 2020), the FIP effect (Laming, 2004, 2015), etc. We are interested in the fact that the ponderomotive force induced by Alfvén waves drives plasma motions along the magnetic field (Lundin and Guglielmi, 2006), and such motions can be simulated and used as an example to compare the effects of using different stratifications.
@@
Figure 8: Longitudinal velocities at $t=1$ s. All settings are listed in Table 1. Except simulation (d), which uses the pHE stratification, all the others use the cHE stratification as the initial condition. Simulations (a) and (b) use assigned constant collisional frequencies. More specifically, (a) is weakly collisional and (b) is strongly collisional. All the other simulations use physical collisional frequencies calculated by Ebysus. Only simulation (e) includes the NEQ ionization/recombination.
The effects of the numerical disturbances appear in the $z$ direction (i.e., the direction of the magnetic field). The Alfvén wave-induced ponderomotive force relates to ${\partial E^{2}_{\perp}}/{\partial z}$, in which $E_{\perp}$ is the electric field perpendicular to the magnetic field (Lundin and Guglielmi, 2006; Martínez-Sykora et al., 2023). Therefore, upward motions (positive longitudinal velocity) are driven by the ponderomotive force, as shown in Figure 8. When it is weakly collisional (panel a), the neutral fluids are decoupled from the ionized fluids, whose motions are driven by the Alfvén wave. Each neutral fluid also moves independently because of the imbalance between gravity and the respective plasma gas pressure, while neither collisions with ionized fluids nor magnetic pressure affect neutral fluids. In contrast, all fluids in the strongly collisional simulation are coupled, behaving as a single-fluid as expected (panel b). More importantly, we can observe that in the strongly collisional simulation, the total velocity is zero before the wave arrives, showing that the cHE stratification is indeed in hydrostatic equilibrium, providing another way to validate the MFMS code – once it is collisionally coupled, the plasma behaves as a single fluid.
@@
The pHE stratification is not further discussed with ionization and recombination due to the numerical instability. We should, however, note that the ionization and recombination timescales are long in the chromosphere (Carlsson and Stein, 2002), and thus the changes in density and temperature would be less drastic when the pHE stratification is used under chromospheric conditions. However, this also means that the ionization fractions (and thus the collisional interactions) in the corresponding simulations would be largely dependent on the initial stratification. We have shown in Section 2, that heavier elements, whose scale heights are different from that of (ionized or neutral) hydrogen, cannot be properly recovered by the pHE stratification, even if the collisional interactions (or other mixing processes) are sufficient to couple the heavy fluids.
@@ Acknowledgements.
FZ thanks Nicolas Poirier, Mats Carlsson, Boris Gudiksen, and Ben Snow for the helpful discussion and Mikolaj Szydlarski for technical support. This research has been supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622, and by computational resources provided by Sigma2, the National Infrastructure for High Performance Computing and Data Storage in Norway. JMS acknowledges support by NASA contracts NNG09FA40C (IRIS) and 80GSFC21C0011 (MUSE), NASA grant 80NSSC26K0018, NSF grants AGS2532363 and AGS2532187. This work used the CHIANTI atomic database. CHIANTI is a collaborative project involving George Mason University, the University of Michigan (USA), the University of Cambridge (UK), and NASA Goddard Space Flight Center (USA). Figures in this work were produced using the Python package Matplotlib (Hunter_2007). Data processing was performed using the NumPy library (Harris et al., 2020) and in-house packages EbysusRunTools_py, atom_py and Helita.
@@ Appendix A Effects of magnetic field
Figure 9: The maximum and averaged $|\mathbf{B}|$ at each height. The maximum value shows a clear inverse power law above $\sim$ 0.1 Mm. The averaged value first drastically increases and then gradually decreases, and exhibits slower variation above $\sim$1 Mm.
@@
Therefore, following the results in Section 3.3, we discuss the effects of adding a magnetic field. The magnetic field in this 2D model is locally concentrated in the photosphere and shows a significant expansion higher up. Consequently, the unsigned mean magnetic field strength decreases over height (Figure 3, bottom). Observing Figure 9, a power law can be seen from the variation of the maximum value at each height, but it is not representative of the whole magnetic field. The averaged value also cannot represent the derivative (i.e., gradient) of the magnetic field. Nevertheless, we do not focus on such details, and the results justify using an expansion assumption to approximate the effect of magnetic pressure.
@@
where $\mathbf{B}_{0}$ and $r_{0}$ are free parameters, to describe the magnetic pressure. With Eq. (15), a smaller $r_{0}$ means that the magnetic field is more concentrated in the photosphere. The pressure gradient can be analytically calculated based on Eq. (15), and directly added to the numerical integration routine. Specifically, $p^{\text{m}}(z_{i})$ should be added at $z_{j}$ in Eq. (11). With a constant $r_{0}$ and varying $\mathbf{B}_{0}$, we have obtained a series of stratifications with the SE ionization, as shown in Figure 10. Note that only the total densities of all fluids are shown for simplicity. Without magnetic pressure (gradient), the cHE stratification has a lower density than the averaged 2D solution. In comparison, adding magnetic pressure gradients significantly alters the solution, as stronger gradients (Figure 11) support higher density. More importantly, the results demonstrate that the current numerical integration routine can be used to calculate cHE stratifications under varying magnetic fields.
Figure 10: cHE stratifications calculated with the given magnetic pressure distributions, $r_{0}=1$ Mm, and SE ionization fractions. The total density when $B_{0}=0$ G is the same as the SE result in Figure 4.
Figure 11: Ratios of the magnetic pressure gradient to the plasma gas pressure gradient. The peaks appear because the plasma gas pressure gradient changes sign. Above the photosphere, the magnetic pressure gradient becomes dominant.
@@
in which $L$ is the radius of the loop. This function is essentially the same as what was used by Karampelas et al. (2019). As $z$ is the height from the footpoint, the gravitational acceleration along the loop axis becomes zero at the loop apex ($z=L$). Plasma along a realistic loop is supported by plasma gas pressure and magnetic pressure/tension, and we can already see that magnetic pressure gradient effectively shapes the 1D stratifications. Particularly, Figure 11 shows that the magnetic pressure description given by Eq. (15) results in dominant magnetic pressure gradients in the upper chromosphere and above. Overall, including Eq. (16) does not significantly change the stratifications here, and the corresponding results are thus not shown.